OpenStudy (anonymous):

What is the value of g[f(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

f(4)=2(4)+1=9 g(9)=2(9)-5=18-5=13

OpenStudy (anonymous):

thankyou!!

OpenStudy (anonymous):

your well